OpenStudy (anonymous):

find the missing value. cos__=2/5 how to do this

OpenStudy (anonymous):

Don't you use the unit circle?

OpenStudy (anonymous):

no idea xD

ganeshie8 (ganeshie8):

you need to use calculator

ganeshie8 (ganeshie8):

\(\cos\_\_ = 2/5 \) \(\_\_= \cos^{-1}(2/5)\)

OpenStudy (usukidoll):

ti-83 calculator xD I think any calculator can do it

ganeshie8 (ganeshie8):

wolfram can do it too http://www.wolframalpha.com/input/?i=cos%5E%28-1%29%282%2F5%29

OpenStudy (usukidoll):

I got 66.42 off of my windows calculator in degree mode

OpenStudy (usukidoll):

yay I got it ! :D

ganeshie8 (ganeshie8):

i think solving analytically is hard we can approximate using newton or some other numerical stuff

OpenStudy (anonymous):

thank you i got it